OpenStudy (anonymous):

What is the solution set of |x – 6| = 8? 14 14 or –14 14 or –2 14 or 2

OpenStudy (nurali):

|x – 6| = 8 x-6=8 or x-6=-8 x=14 or x=-2

OpenStudy (anonymous):

my first try I gor x=14 or x=-2 was I doing it right